My son needs your help in solving this problem.
Find the solution set for the system of equations
3X + 2Y = 3
-5X + Y = 8
Thanks

Answer by Earlsdon(6294)

Solve the system of equations:
1) 3x + 2y = 3
2) -5x + y = 8 Multiply equation 2) by 2 then subtract equation 1) from the result and solve for x.
2) -10y + 2y = 16
1) -(3x + 2y = 3)
-----------------
3) -13x = 13 Divide both sides by -13
x = -1 Now substitute this into equation 1) and solve for y.
1) 3(-1) + 2y = 3
-3 + 2y = 3 Add 3 to both sides.
2y = 6 Divide both sides by 2.
y = 3
Solution: (-1, 3)